OpenStudy (anonymous):

Rationalize the denominator in the expression. sq. root of 245x^7 over sq. root of y^9

OpenStudy (campbell_st):

\[\frac{\sqrt{245x^7}}{\sqrt{y^9}} \times \frac{\sqrt{y^9}}{\sqrt{y^9}} = \frac{\sqrt{245x^7} \times \sqrt{y^9}}{\sqrt{y^9}\times \sqrt{y^9}}\] you should be able to finish it from here